OpenStudy (anonymous):

solve each system by substitution. -8x-4y=-4 y=x-8

OpenStudy (compassionate):

-8x - 4y = - 4 Now you want to plug in y, we know y = x - 8 -8x - 4(x - 8) = -4 Now solve for x -8x - 4x +32 = -4 Now add like terms -12x + 32 = -4 Subtract 32 -12x = -36 Divide by -12 \[x = \frac{ -12 }{ -36 }\] x = 3 Now take that 3 and plug it into your equation for x -8(3) - 4y = -4 Now you just need to solve for y Your x coordinate is (3, ?) You'll get your y coordinate by solving for y

OpenStudy (anonymous):

Multiply the second equation by 4 so that your coefficients of y are (-) of each other: 4y=4x-32 Then add the two equations. Since your y coeeficients were opposite (-) they will cancel out and go to 0, leaving you with one variable: -8x=-4+4x-32 Subtract 4x, and combine your basic numbers -12x=-36 Divide by -12 x=3 Now plug this into either equation (second is simpler) to get y: y=3-8=-5 So the solution is (3,-5)

OpenStudy (anonymous):

can you show me how to plug this into the equation